OpenStudy (anonymous):

Solve the system of linear equations below. y = -2x + 19 y = x + 7

OpenStudy (anonymous):

i can

OpenStudy (anonymous):

solve for u

OpenStudy (anonymous):

ok i can help

OpenStudy (anonymous):

omg i would love that soo much

OpenStudy (anonymous):

love what exactly are u looking for

OpenStudy (anonymous):

they're asking me to solve for (x) and (y) but im not good with math i dont understand here are my options x = 4, y = 11 x = -12, y = -5 x = -4, y = 3 x = 8, y = 3

OpenStudy (anonymous):

if am not wrong i think for the second one it -7,+7 or 7,-7 i forgot the order haha

OpenStudy (anonymous):

or if i think that the answer i got

OpenStudy (anonymous):

i dont see that in any of my options

OpenStudy (acxbox22):

since bot equations are equal to y set them equal to each other -2x+19=x+7 now solve for x and tell me what you get

OpenStudy (campbell_st):

since both equations equal y, you can equate them to solve for x y = -2x + 19 y = x + 7 then x + 7 = -2x + 19 start by adding 2x to both sides of the equation 3x + 7 = 19 now solve for x. when you get x, substitute it into either equation to find y. hope it helps

OpenStudy (anonymous):

still confused so you can just put any number in ?

OpenStudy (campbell_st):

so you can't solve 3x + 7 = 19 is that correct..?

OpenStudy (anonymous):

i can slove it (X)= 4

OpenStudy (acxbox22):

yes now we have x so you can substitute 4 for x in either equation and solve for y

OpenStudy (acxbox22):

so y=4+7 y=??

OpenStudy (anonymous):

y=11 wow thank both of you soo much that helped alot @campbell_st @acxbox22